# Closed-Form Training Dynamics Reveal Learned Features and Linear Structure in Word2Vec-like Models

### v2

@@ Appendix B Additional figures
Figure 5: word2vec eigen-features align very closely with QWEMs’ and less with PPMI. The heatmap indicates the squared overlaps between the latent feature vectors (i.e., the left singular vectors) between the three models considered. This is a quantitative version of the table in Figure 3.
Figure 6: The empirical observations regarding linear representations extend across the semantic classes in the analogy dataset. We show that across analogy categories, the corresponding task vectors exhibit the geometric structure discussed in Figure 4.
@@ Appendix C Proofs
Proof. Define ${\bm{M}}\coloneqq{\bm{W}}{{\bm{W}}}^{\top}$. Rewriting Equation 3 and plugging in Equation 5 and 1,
@@
[…]
Proof. The formulation as a weighted matrix factorization follows from Equation 30. In the case that ${\bm{G}}$ is rank 1, substituting in ${\bm{\Gamma}}$, we find
@@
[…]
@@ D.1 Relation to PMI
where the function $\Delta(x)$ yields the fractional deviation away from independent word statistics, in terms of some small parameter $x$ of our choosing (so that $\Delta(0)=0$). This setup allows us to Taylor expand quantities of interest around $x=0$. A judicious choice will produce terms that cancel the $-\frac{1}{2}\Delta^{2}$ that arises from the Taylor expansion of $\log(1+\Delta)$, leaving only third-order corrections. One such example is $\Delta(x)=2x/(2-x)$, which yields
@@
This calculation reveals that ${\bm{M}}^{\star}$ learns a very close approximation to the PMI matrix; the leading correction is third order. However, ${\bm{M}}^{\star}$ is much friendlier to least-squares approximation, since $x$ is bounded ($-2\leq{\bm{M}}^{\star}_{ij}\leq 2$).
@@ D.2 Relation to SimCLR
If the model is in a linearized regime, we may approximate ${\bm{S}}_{\theta}(i,j)\approx{{\bm{f}}}^{\top}_{i}{\bm{f}}_{j}$ for some linearized feature vectors ${\bm{f}}$. Then the loss can be written as an unweighted matrix factorization problem using exactly the same argument as in 1. Thus, we expect that vision models trained under the SimCLR loss in a linearized regime will undergo stepwise sigmoidal dynamics. This provides an explanation for the previously unresolved observation in Simon et al. (2023b) that vision models trained with SimCLR from small initialization exhibit stepwise learning.
@@
The current LLM paradigm demonstrates that the next-token distribution is largely sufficient for most downstream tasks of interest. The next-token prediction (NTP) task aims to model the probability of finding word $i$ given a preceding window of context tokens of length $L-1$. Therefore, the NTP target is an order-$L$ tensor that captures the joint distribution of length-$L$ contexts. NTP thus generalizes the word embedding task. Both QWEM and LLMs are underparameterized models that learn internal representations with interpretable and task-relevant vector structure. Both are trained using self-supervised gradient descent algorithms, implicitly learning a compression of natural language statistics by iterating through the corpus.

### v3

@@ Appendix B Additional figures
Figure 5: word2vec eigen-features align very closely with QWEMs’ and less with PPMI. The heatmap indicates the squared overlaps between the latent feature vectors (i.e., the left singular vectors) between the three models considered. This is a quantitative version of the table in Figure˜3.
Figure 6: The empirical observations regarding linear representations extend across the semantic classes in the analogy dataset. We show that across analogy categories, the corresponding task vectors exhibit the geometric structure discussed in Figure˜4.
Figure 7: Ablation tests. Starting from vanilla word2vec (top left), we separately add in both the quartic approximation (top right) and the reweighting condition ˜1 (bottom left). We include QWEM on the bottom right (i.e., both effects). The quartic approximation hardly changes the singular value dynamics at all; the clean mode separation is due to the reweighting. However, as is common with logistic losses, the weights begin to diverge at late times. Using the quartic loss prevents this. The experiments here use a 41-million-token mixture of the Corpus of Contemporary American English and the News on the Web dataset. We use a vocabulary size of 20,000; the embedding dimension is 150; 2 negative samples per positive sample; and a context length of 16.
Figure 8: Although the analogy geometry degrades with $d$, relative scores remain high. We plot two different measures of analogy geometry across $d$ to illustrate subtleties associated with the top-1 accuracy metric. For an analogy $a:b::c:w$, we denote $\hat{\bm{x}}:={\bm{c}}+({\bm{b}}-{\bm{a}})$ the prediction for the fourth word. We plot the average cosine similarity between $\hat{\bm{x}}$ and the true $\bm{w}$} over the analogy family. The prediction degrades dramatically as the embedding dimension increases, complementing the observation in Figure˜4 that the SNR decreases with $d$. However, this geometric breakdown fails to capture the fact that all embedding vectors separate as $d$ increases; if we normalize by the maximum cosine similarity between $\bm{w}$} and non-analogy embeddings, the score remains roughly stable at large $d$. This provides an explanation for why top-1 accuracy often remains high despite the breakdown of geometric analogical structure.
@@ Appendix C Proofs
Proof. Define ${\bm{M}}\coloneqq{\bm{W}}{{\bm{W}}}^{\top}$. Rewriting Equation˜3 and plugging in Equation˜5 and ˜1,
@@
Since $P_{ij}$, $P_{i}P_{j}$, $\Psi^{+}_{ij}$, $\Psi^{-}_{ij}$ are all real symmetric, so is $\bm{M}$$^{\star}$, so it has an eigendecomposition. By the Eckart-Young-Mirsky theorem, the loss-minimizing $\bm{M}$} must be the truncated SVD ${\bm{M}}^{\star}_{[d]}$, whose symmetric factors are exactly given by Equation˜7. $\qquad\blacksquare$
@@
Proof. The formulation as a weighted matrix factorization follows from Equation˜30. In the case that $\bm{G}$} is rank 1, substituting in $\bm{\Gamma}$}, we find
@@
whose symmetric factors are exactly given by Equation˜9. $\qquad\blacksquare$
@@ D.1 Relation to PMI
where the function $\Delta(x)$ yields the fractional deviation away from independent word statistics, in terms of some small parameter $x$ of our choosing (so that $\Delta(0)=0$). This setup allows us to Taylor expand quantities of interest around $x=0$. A judicious choice will produce terms that cancel the $\frac{1}{2}\Delta^{2}$ that arises from the Taylor expansion of $\log(1+\Delta)$, leaving only third-order corrections. One such example is $\Delta(x)=2x/(2-x)$, which yields
@@
This calculation reveals that $\bm{M}$$^{\star}$ learns a very close approximation to the PMI matrix; the leading correction is third order. However, $\bm{M}$$^{\star}$ is much friendlier to least-squares approximation, since $x$ is bounded ($2\leq{\bm{M}}^{\star}_{ij}\leq 2$).
@@ D.2 Relation to SimCLR
If the model is in a linearized regime, we may approximate ${\bm{S}}_{\theta}(i,j)\approx{{\bm{f}}}^{\top}_{i}{\bm{f}}_{j}$ for some linearized feature vectors $\bm{f}$}. Then the loss can be written as an unweighted matrix factorization problem using exactly the same argument as in ˜1. Thus, we expect that vision models trained under the SimCLR loss in a linearized regime will undergo stepwise sigmoidal dynamics. This provides an explanation for the previously unresolved observation in Simon et al. (2023b) that vision models trained with SimCLR from small initialization exhibit stepwise learning.
@@
The current LLM paradigm demonstrates that the next-token distribution is largely sufficient for most downstream tasks of interest. The next-token prediction (NTP) task aims to model the probability of finding word $i$ given a preceding window of context tokens of length L1L-1. Therefore, the NTP target is an order-$L$ tensor that captures the joint distribution of length-$L$ contexts. NTP thus generalizes the word embedding task. Both QWEM and LLMs are underparameterized models that learn internal representations with interpretable and task-relevant vector structure. Both are trained using self-supervised gradient descent algorithms, implicitly learning a compression of natural language statistics by iterating through the corpus.
@@
D.4 Relation to neural tangent kernel.
Our result describing an implicit bias towards low rank directly contrasts the well-studied neural tangent kernel training (NTK) regime (Jacot et al., 2018; Chizat et al., 2019; Karkada, 2024). Here we compare the two learning regimes.
Learning with NTK.
  • Learning dynamics and generalization performance can be solved (Lee et al., 2019; Bordelon et al., 2020; Simon et al., 2023a).
  • Extreme over-parameterization is required; large finite-width corrections at practical widths (Huang & Yau, 2020).
  • Model weights do not learn task-relevant features.
  • Optimization remains in a locally convex region of the loss landscape (Chizat et al., 2019).
Learning from small initialization.
  • Learning dynamics are generally complicated; can be solved in very simple cases (linear networks, special data distributions, etc.).
  • Behavior is qualitatively consistent across network widths, with only moderate finite-width corrections (Vyas et al., 2023).
  • Model weights learn task-relevant features.
  • Optimization tends to pass near a sequence of saddle points. (Baldi & Hornik, 1989; Jacot et al., 2021).